I divided 1.37 by 1.09 and am not getting the answer which is 12.01%. I know i have a math error can you please write out the steps how they got this [1+z3]3= [1+z2]? [1 +f12) (1 + 0.11)3= (1 + 0.09)1 [1 + $1,2)”; $1,2 = 12.01%
Answer #1

I think you may be missing to square root the answer.

(1+0.11)3 = (1+0.09)1 x [1 + f1,2]2

1.37 = 1.09 x [1 + f1,2]2

1.2547 = [1 + f1,2]2

Square Root of (1.2547) = 1 + f1,2

1.1201 = 1 + f1,2

f1,2 = 1.1201 - 1

f1,2 = 12.01%